OpenStudy (anonymous):

need help can anyone help me with regard of circles i mean the needed formulas in solving for them ...\ im just reviewing for our exam in analytic geometry please?

OpenStudy (agent0smith):

Like this? \[\large (x-h)^2 +(y-k)^2 = r^2\] general formula for a circle, centred at (h, k) with radius r.

OpenStudy (agent0smith):

\[\large Area = \pi r^2\] \[\large Circumference = 2 \pi r\]

OpenStudy (agent0smith):

arc length L, circumference C, theta is angle \[\large \frac{ \theta }{ 360} = \frac{ L }{ C }\]

OpenStudy (anonymous):

yeah yeah tnx

OpenStudy (anonymous):

do you know how the normal form of a line works? i mean how to do it?

OpenStudy (aravindg):

this would surely help http://www.schooltutoring.com/help/equation-of-a-line-normal-form/

OpenStudy (anonymous):

i want to thank the both of you :) i passed :)